OpenStudy (anonymous):

how do you write 84^-3/4 in radical notation

OpenStudy (anonymous):

84=2*2*3*7 so it becomes\[\sqrt[-3/4]{21} \sqrt{2} \] so \[\sqrt{2}/\sqrt[4]{21^{3}}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (unklerhaukus):

\[84^{-3/4}\\ =(2^2\cdot3\cdot7)^{-3/4}\\=\frac1{(2^2\cdot3\cdot7)^{3/4}}\\ =\frac1{\sqrt[4]{(2^2\cdot3\cdot7)^{3}}}\\ =\frac1{\sqrt[4]{2^6\cdot3^3\cdot7^3}}\\ =\frac1{\sqrt[4]{2^4\cdot2^2\cdot3^3\cdot7^3}}\\ =\frac1{2\sqrt[4]{\cdot2^2\cdot3^3\cdot7^3}}\]